the solution of the given linear differential equation x^2 y'+x(x+2)=e^x for x>0


1
Expert's answer
2021-11-12T12:32:32-0500

"\begin{gathered}
x^{2} y^{\prime}+x(x+2)=e^{x} \\
x^{2} y^{\prime}=e^{x}-x(x+2) \\
y^{\prime}=\frac{e^{x}-x^{2}-2 x}{x^{2}} \\
\frac{\mathrm{d} y}{\mathrm{~d} x}=\frac{e^{x}-x^{2}-2 x}{x^{2}} \\
\mathrm{~d} y=\frac{\left(e^{x}-x^{2}-2 x\right) \mathrm{d} x}{x^{2}} \\
\int 1 \mathrm{~d} y=\int \frac{e^{x}}{x^{2}}-\frac{2}{x}-1 \mathrm{~d} x \\
y=-2 \ln (x)+\mathbf{E i}(x)-\frac{e^{x}}{x}-x+C
\end{gathered}"